OpenStudy (kitkat1):

Medal for best answer!! Simplify square root of 3 multiplied by the fifth root of 3 . 3 to the power of 1 over 10 3 to the power of 3 over 5 3 to the power of 3 over 5 3 to the power of 7 over 10

OpenStudy (alexandervonhumboldt2):

ok so look: lets convert to exponent form: sqrt3=3^1/2 and 5throot(3)=3^1/5

OpenStudy (alexandervonhumboldt2):

3^1/2 * 3^1/5=3^(1/2+1/5)=3^7/10

OpenStudy (kitkat1):

so its D?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (kitkat1):

Can u help me with other one?